Grade 12th passPhysical Chemistry

Calculate the total pressure in a mixture of 3.5 g of dinitrogen and 16g of dioxygen confined in a vessel of 2dm3 at270°C.(R=0.083 bar dm3 K-1 mol-1)​

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer:

Given,

Mass of oxygen = 8 g, molar mass of oxygen = 32 g/mol

Mass of hydrogen = 4 g, molar mass of hydrogen = 2 g/mol

Therefore amount of oxygen = 8/ 32 = 0.25 mol

And amount of hydrogen = 4/2 = 2 mol

From the gas equation PV = n RT, we get,

P X 1 = (0.25 + 2) X 0.083 X 300 = 56.02 bar

Hence, the total pressure of the mixture is 56.02 bar

Thanks